### v1

# Adaptive Teaming in Multi-Drone Pursuit: Simulation, Training, and Deployment

Abstract
Adaptive teaming, the ability to collaborate with unseen teammates without prior coordination, remains an underexplored challenge in multi-robot collaboration. This paper focuses on adaptive teaming in multi-drone cooperative pursuit, a critical task with real-world applications such as border surveillance, search-and-rescue, and counter-terrorism. We first define and formalize the Adaptive Teaming in Multi-Drone Pursuit (AT-MDP) problem and introduce AT-MDP framework, a comprehensive framework that integrates simulation, algorithm training and real-world deployment. AT-MDP framework provides a flexible experiment configurator and interface for simulation, a distributed training framework with an extensive algorithm zoo (including two newly proposed baseline methods) and an unseen drone zoo for evaluating adaptive teaming, as well as a real-world deployment system that utilizes edge computing and Crazyflie drones. To the best of our knowledge, AT-MDP framework is the first adaptive framework for continuous-action decision-making in complex real-world drone tasks, enabling multiple drones to coordinate effectively with unseen teammates. Extensive experiments in four multi-drone pursuit environments of increasing difficulty confirm the effectiveness of AT-MDP framework, while real-world deployments further validate its feasibility in physical systems. Videos and code are available at https://sites.google.com/view/at-mdp.
Machine Learning, ICML
1 Introduction
Multi-drone pursuit is an increasingly critical task with wide-ranging real-world applications, such as disaster response, border surveillance, search-and-rescue operations, and many more (Chung et al., 2011; Zhang et al., 2023; Queralta et al., 2020). All of these scenarios are highly dependent on the ability of autonomous drones to coordinate effectively in dynamic environments to pursue and track uncooperative targets. Most existing approaches rely on pre-coordinated strategies, where agents either follow predefined coordination mechanisms, such as conventions, role assignments, and communication protocols, or learn them through extensive interaction over time (Rahman et al., 2024). These methods generally fall into two categories: traditional optimization-based approaches (Shah & Schwager, 2019; Janosov et al., 2017b; Zhou et al., 2016),and reinforcement learning-based approaches (Zhang et al., 2023; Chen et al., 2024; Qi et al., 2024; de Souza et al., 2021; Li et al., 2019; Matignon et al., 2007).
Figure 1: Pre-coordinated teaming: Agents coordinate using predefined mechanisms—such as conventions, role assignment protocols, and communication strategies—or by learning these mechanisms through extensive interaction over time (Rahman et al., 2024). Adaptive teaming: The assumed convention or mechanism is disrupted or fails, requiring adaptation to unseen teammates without prior coordination methods.
However, real-world scenarios are inherently uncertain, and unpredictable, often necessitating rapid adaptation to evolving team compositions. Instead of relying on pre-coordinated teaming, drones must seamlessly integrate with unforeseen teammates in real time. For example, in disaster response, damaged drones may need rapid replacement to maintain search coverage, while in border surveillance, specialized drones might temporarily join to intercept high-priority targets. Adaptive teaming is capable of addressing this critical gap by enabling drones to dynamically adjust their coordination strategies with unforeseen teammates, without relying on prior coordination, as shown in Fig. 1.
To address the challenge of working with unseen drone teammates, we first define and formalize the Adaptive Teaming in Multi-Drone Pursuit (AT-MDP) problem and propose the AT-MDP framework, a comprehensive framework that seamlessly integrates simulation, training, and real-world deployment. Specifically, the simulation component enables flexible customization of multi-drone pursuit scenarios through an environment configurator and provides a realistic simulated environment for training and evaluation via the environment interface. The training component leverages a distributed framework for algorithm learning, incorporating a comprehensive algorithm zoo that includes two newly proposed baseline methods designed for adaptive teaming, along with an unseen drone zoo to assess generalization to diverse teammate behaviors. Finally, the deployment component bridges simulation to real-world applications by integrating a motion capture system and edge nodes, enabling real-time data exchange and decision-making for physical drone coordination.
AT-MDP framework also advances adaptive teaming in machine learning by enabling multiple drones to dynamically collaborate with unseen teammates in complex, continuous-action environments, pushing the boundaries of multi-agent coordination in real-world scenarios. In machine learning, zero-shot coordination (ZSC) (Hu et al., 2020) and ad-hoc teamwork (AHT) (Stone et al., 2010) offer alternative paradigms for handling unseen partners. However, most of the research in this area is confined to training a single learner in simulated 2D video games with discrete action spaces, such as Overcooked (Li et al., 2023; Wang et al., 2024d), Hanabi (Hu et al., 2020; Lucas & Allen, 2022; Canaan et al., 2022; Bard et al., 2020), and Predator-Prey (Barrett et al., 2011; Papoudakis et al., 2021). A recent advancement, NAHT (Wang et al., 2024a), extends AHT to control multiple learners interacting with multiple unseen partners. However, it remains restricted to the SMAC benchmark based on video games (Samvelyan et al., 2019) with discrete actions, limiting its applicability to real-world continuous-action drone tasks.
To validate AT-MDP framework, we conducted extensive experiments in four multi-drone pursuit environments of increasing difficulty, evaluating their adaptability to diverse unseen teammate dynamics. The experiments incorporate multiple unseen drone protocols to ensure a broad spectrum of coordination challenges. The results show that our proposed baseline methods effectively address AT-MDP problem, consistently outperforming existing approaches in coordination success and adaptability. Furthermore, real-world deployment on Crazyflie drones demonstrates the feasibility of AT-MDP framework in physical systems, while highlighting opportunities for further improvements in algorithms, more realistic task scenarios, and enhanced sim-to-real transfer. A case study in a high-complexity scenario further highlights the emergence of adaptive strategies in dynamic, real-world-inspired conditions.
The main contributions of this paper are threefold: (1) To the best of our knowledge, this work is the first to formally define and formulate the adaptive teaming problem in multi-drone pursuit, advancing multi-agent coordination in complex, continuous-action real-world scenarios. (2) We propose a comprehensive AT-MDP framework that integrates flexible simulation, algorithm training, and real-world deployment, paving the way for further research in this area; (3) We validate the effectiveness and feasibility of our baseline methods through extensive simulations and real-world experiments with diverse, previously unseen drone partners.
Table 1: Comparison of AT-MDP framework with main related works. Grey rows represent literature on multi-drone pursuit, while pink rows highlight adaptive teaming studies from the machine learning field. “AT w/o TM” and “AT w/ TM’ denotes adaptive teaming without and with, respectively. A complete tabular comparison of related works is provided in Table 4 in the Appendix.
Related Work Problem Setting Task Method
# Learner # Unseen # Evader Action Space Main Related Task Real-world? AT w/o TM? AT w/ TM?
Voronoi Partitions (Zhou et al., 2016) Multi 0 1 Continuous Pursuit–evasion Game No No No
Bio-pursuit (Janosov et al., 2017b) Multi 0 Multi Continuous Prey–predator Game No No No
M3DDPG (Li et al., 2019) Multi 0 1 Continuous Prey–predator Game No No No
DualCL (Chen et al., 2024) Multi 0 1 Continuous Multi-drone Pursuit No No No
Pursuit-TD3 (de Souza et al., 2021) Multi 0 1 Continuous Multi-drone Pursuit Yes No No
DACOOP-A (Zhang et al., 2023) Multi 0 1 Discrete Multi-drone Pursuit Yes No No
HOLA-Drone (Li et al., 2024a) 1 Multi Multi Continuous Multi-drone Pursuit Yes Yes No
Other-play (Hu et al., 2020) 1 1 0 Discrete Lever Game; Hanabi No Yes No
COLE (Li et al., 2024b) 1 1 0 Discrete Overcooked No Yes No
PLASTIC (Barrett et al., 2017) 1 Multi Multi Discrete Prey-predator Game No No Yes
AATeam (Chen et al., 2020) 1 1 2 Discrete Half Field Offense No No Yes
CIAO (Wang et al., 2024b) 1 Multi Multi Discrete LBF; Wolfpack No No Yes
NAHT (Wang et al., 2024a) Multi Multi Multi Discrete StarCraft2 No No Yes
Our ATMDP Multi Multi Multi Continuous Multi-drone Pursuit Yes Yes Yes
2 Problem Formulation and Related Work
In this section, we first formalize the concept of adaptive teaming in multi-drone pursuit. Next, we discuss the relationship between our work and prior research, as summarized in Table 1. A more detailed literature review and complete tabular comparison of related works can be found in the Appendix A.
2.1 Problem Formulation
Definition 2.1 (Adaptive Teaming in Multi-Drone Pursuit).
Adaptive teaming in multi-drone pursuit involves training a set of $N\in\{1,2,\dots\}$ drone agents, referred to as learners, to dynamically coordinate with $M\in\{1,2,\dots\}$ previously unseen partners. The objective is to pursue $K\in\{1,2,\dots\}$ targets without collisions, optimizing the overall return.
Let ${\mathcal{C}}$ represent the cooperative team, comprising $N$ learners and $M$ uncontrolled teammates. The set of uncontrolled teammates is denoted by ${\mathcal{U}}$. In the multi-drone pursuit task, there exists a set of opponents, denoted as ${\mathcal{E}}$.
Adaptive teaming can be effectively modeled as an extended Adaptive Teaming Decentralized Partially Observable Markov Decision Process (AT-Dec-POMDP). AT-Dec-POMDP is defined by the tuple $({\mathcal{S}},{\mathcal{C}},{\mathcal{A}},{\mathcal{P}},{\mathcal{P}}_{\text{u}},r,{\mathcal{O}},\gamma,T)$, where: where ${\mathcal{S}}$ is the joint state space; ${\mathcal{C}}$ is the set of cooperative agents, including learners (${\mathcal{N}}$) and uncontrolled teammates (${\mathcal{M}}$); ${\mathcal{A}}=\times_{j=1}^{C}{\mathcal{A}}^{j}$ is the joint action space, where $C=N+M$ is the team size; ${\mathcal{P}}_{\text{u}}({\mathcal{M}}|{\mathcal{U}})$ is the uncontrolled teammate sampling function, which defines the probability of sampling a subset ${\mathcal{M}}\subseteq{\mathcal{U}}$ of size $M$ from the set of all uncontrolled teammates ${\mathcal{U}}$; ${\mathcal{P}}(s’|s,a)$ is the transition probability function, representing the probability of transitioning to state $s’\in{\mathcal{S}}$ given the current state $s\in{\mathcal{S}}$ and joint action $a\in{\mathcal{A}}$; $r(s,a)$ is the reward function, representing the team’s reward in state $s$ after taking action $a$; ${\mathcal{O}}$ is the joint observation space, with ${\mathcal{O}}(o|s)$ describing the probability of generating observation $o$ given state $s$; $\gamma\in[0,1]$ is the discount factor; and $T$ is the task horizon.
At each time step $t$, the AT-Dec-POMDP is in state $s_{t}\in{\mathcal{S}}$ and generates a joint observation $o_{t}=(o_{t}^{1},\dots,o_{t}^{C})\sim{\mathcal{O}}(\cdot|s_{t})$. Each agent $j\in{\mathcal{C}}$ observes $o_{t}^{j}$ and maintains its own action-observation trajectory $\tau_{t}^{j}=(o_{0}^{j},a_{0}^{j},\dots,o_{t-1}^{j},a_{t-1}^{j},o_{t}^{j})$. Furthermore, we denote ${\mathcal{T}}^{j}$ as the set of all possible partial observations and action histories for the agent $j$. Using its policy $\pi^{j}$, the agent selects an action $a_{t}^{j}\in{\mathcal{A}}^{j}$. The policies of the $N$ learners, denoted by $\pi^{i}$ for $i\in{\mathcal{N}}$, are learnable.
We define the policy in two approaches: with and without teammate modeling. Adaptive teaming without teammate modeling is closely related to the zero-shot coordination problem (Hu et al., 2020; Carroll et al., 2019), where learners must rapidly coordinate with unseen teammates. Specifically, the policy for a learner $i$ is represented as $\pi^{i}(a_{t}^{i}\mid\tau_{t}^{i})$, where $\tau_{t}^{i}$ denotes the learner’s observation history up to time $t$. On the other hand, adaptive teaming with teammate modeling aligns closely with the ad-hoc teamwork paradigm (Stone et al., 2010), where learners explicitly model their teammates using a separate teammate model $f$. In this case, the policy is defined as $\pi^{i}(a_{t}^{i}\mid\tau_{t}^{i},f(\tau_{t}^{i}))$. The joint action $a_{t}=(a_{t}^{1},\dots,a_{t}^{C})$ determines the next state $s_{t+1}\sim{\mathcal{P}}(s_{t+1}\mid s_{t},a_{t})$, and all agents receive a shared reward $r(s_{t},a_{t})$.
The goal of adaptive teaming is to learn policies $\{\pi^{i}\}_{i\in{\mathcal{N}}}$ that maximize the expected discounted return: ${\mathcal{J}}=\mathbb{E}[R(\tau)]=\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}r(s_{t},a_{t})\right],$ where $\tau$ denotes the joint trajectory.
Figure 2: Overview of AT-MDP framework. AT-MDP framework consists of three core components: Simulation, Training, and Deployment. The Simulation module provides an Environment Configurator and Interface for flexible scenario customization and a simulated training environment. The Training module operates in a distributed framework, featuring an Algorithm Zoo and an Unseen Drone Zoo for evaluating adaptive teaming. The Deployment module integrates learned policies into real-world tasks using a motion capture system and edge nodes for real-time decision-making, seamlessly bridging simulation, training, and deployment.
2.2 The Relation to Prior Works
In this section, we highlight the key distinctions between AT-MDP framework and prior works. Tables 1 and Table 4 provide a comprehensive comparison across various dimensions, including problem formulation, scope of tasks, and methodological approach. AT-MDP framework introduces a novel adaptive teaming framework tailored for multi-drone pursuit, setting it apart from existing methods. By pushing the boundaries of existing research, AT-MDP framework paves the way for breakthroughs in multi-drone applications for disaster response, border surveillance, and counter-terrorism.
Comparison with Multi-Drone Pursuit Literature (Gray Rows). The studies highlighted in gray focus on pursuit-evasion tasks, with several works (Zhou et al., 2016; Janosov et al., 2017b; Li et al., 2019; Chen et al., 2024) restricted to simulated environments without real-world deployment. Others, such as Pursuit-TD3 (de Souza et al., 2021) and DACOOP-A (Zhang et al., 2023), operate under pre-coordinated settings, where all pursuers are jointly trained to optimize performance in real-world drone pursuit tasks. Unlike these approaches, while HOLA-Drone (Li et al., 2024a) is a recent attempt at adaptive teaming, it is limited to a single learner coordinating with multiple teammates without explicit teammate modeling. AT-MDP framework overcomes these limitations by enabling multiple learners to dynamically adapt and coordinate with unseen teammates in continuous-action real-world multi-drone pursuit tasks.
Comparison with Adaptive Teaming Literature (Pink Rows). In the machine learning community, zero-shot coordination (ZSC) and ad-hoc teamwork (AHT) are closely related to adaptive teaming but have been explored primarily in discrete-action video game environments, as shown in the pink rows of Table 1. Almost all ZSC methods, including Other-play (Carroll et al., 2020), TrajDi (Lupu et al., 2021), MEP (Zhao et al., 2023), LIPO (Charakorn et al., 2023), COLE (Li et al., 2024b), and ZSC-Eval (Wang et al., 2024d), adopt a two-player game formulation with discrete action spaces. For AHT, most existing methods can be interpreted as a specific case of the AT-Dec-POMDP formulation with $N=1$, where a single learner adapts to unseen teammates (Stone et al., 2010; Rahman et al., 2022, 2024; Barrett et al., 2017; Chen et al., 2020; Wang et al., 2024c). The recent NAHT (Wang et al., 2024a) extends AHT to a multi-learner setting, closely resembling the AT-Dec-POMDP formulation with teammate modeling. However, NAHT remains limited to discrete-action decision-making in video game environments, whereas AT-MDP framework advances adaptive teaming by enabling continuous-action decision-making in real-world multi-drone pursuit tasks.
3 The Framework: AT-MDP
To address the adaptive teaming problem in the multi-drone pursuit task, we propose a comprehensive framework called AT-MDP framework, encompassing three key components: Simulation, Training, and Deployment, as illustrated in Fig. 2. The Simulation component allows flexible customization of multi-drone pursuit environments through the environment configurator, while the environment interface provides a simulated environment for training and evaluation. The Training component leverages a distributed training framework to train algorithms through interaction with the environment interface within the simulation. It also includes an algorithm zoo that features a series of baseline methods for solving AT-MDP problem, as well as an unseen drone zoo that serves as a repository of various unseen drone policies to test adaptive teaming capabilities. Finally, the Deployment component facilitates real-world applications using a motion capture system and edge nodes, enabling real-time data exchange and decision-making. Section 3.2 introduces the simulation and deployment modules in detail, while Section 3.1 provides a detailed explanation of the training component.
3.1 Training
The training module comprises three key components: a distributed training framework, an algorithm zoo, and an unseen drone zoo. To enable efficient interaction with environments, AT-MDP framework supports distributed environment sampling, ensuring scalability and adaptability in training processes. The algorithm zoo and the unseen drone zoo are the core components of the training module, providing a series of baseline methods for AT-MDP problem and a diverse population of drone policies for evaluation.
Algorithm Pool. The algorithm pool in AT-MDP framework includes MARL-based pre-coordinated approaches, self-play frameworks, population-based training (PBT) strategies, and our proposed methods for adaptive teaming. For MARL-based approaches, we use the MAPPO algorithm (Yu et al., 2022), a strong baseline designed for multi-drone pursuit tasks, which leverages centralized training with decentralized execution for effective coordination. Self-play and PBT (Carroll et al., 2019) serve as widely used methods for zero-shot coordination, with self-play enabling agents to iteratively learn through interactions with copies of themselves and PBT allowing a population of models to explore various strategies and share knowledge for improved generalization. Both self-play and PBT are implemented using the Independent PPO (Schulman et al., 2017) framework, ensuring scalable and robust training for adaptive teaming tasks.
In addition to existing baseline methods, we propose a novel Open-Ended Population Training (OPT) algorithm to better address AT-MDP problem without teammate modeling. It begins by defining a population of drone strategies, $\Pi=\{\pi_{1},\pi_{2},\cdots,\pi_{n}\}$, where interactions among $C$ teammates are represented as a hypergraph $\mathcal{G}=(\Pi,\mathcal{E},\mathbf{w})$. Here, $\Pi$ represents the set of strategies, $\mathcal{E}$ denotes the interaction relationships, and $\mathbf{w}$ captures the results of these interactions. To evaluate coordination ability, the preference hypergraph $\mathcal{P}\mathcal{G}$ is derived from $\mathcal{G}$, retaining only hyperedges that represent optimal teammate configurations. The cooperative ability of each strategy is quantified by its preference centrality: $\eta_{\Pi}(i)=\frac{d_{\mathcal{P}\mathcal{G}}(i)}{d_{\mathcal{G}}(i)},$ where $d_{\mathcal{P}\mathcal{G}}(i)$ and $d_{\mathcal{G}}(i)$ are the incoming degrees in $\mathcal{P}\mathcal{G}$ and $\mathcal{G}$, respectively. To enhance adaptability, the max-min preference oracle alternates between two steps: identifying the most challenging teammate subsets (minimization) and updating the learner set to maximize performance against these subsets (maximization). The oracle updates the learner set by solving: $\mathcal{N}^{\prime}=\arg\max_{\mathcal{N}}\min_{\mathcal{M}\subseteq\Pi_{-\mathcal{N}}}\phi_{\mathcal{M}}(\mathcal{N}),$ where $\mathcal{N}$ is the learners, $\Pi_{-{\mathcal{N}}}$ the objective $\phi_{\mathcal{M}}(\cdot)$ is formulated using the extended curry operator (Balduzzi et al., 2019), which was originally developed for two-player games. The function $\phi$ evaluates the interactions between $N$ learners and $M$ teammates, producing a real-valued performance outcome. The max-min preference oracle forms the foundation of the OPT algorithm, which evolves the strategy population $\Pi$ and iteratively refines the learner set $\mathcal{N}$ over generations. By dynamically recalibrating the training objectives, OPT promotes robust coordination with diverse, unseen teammates, ensuring adaptability in complex real-world scenarios. For detailed algorithmic steps and implementation, please refer to the Appendix B.
We propose an adaptive teaming with modeling (ATM) baseline to efficiently model unseen drone teammates. ATM incorporates an additional teammate modeling network built on MAPPO (Yu et al., 2022). The modeling network, denoted as $f$, generates team encoding vectors that characterize the strategies of teammates. Specifically, the teammate modeling network $f_{\text{tm}}:{\mathcal{T}}^{i}\rightarrow\mathbb{R}^{n}$, parameterized by $\theta_{tm}$, processes the modeling agent’s history of local observations and actions. The resulting team encoding vector is integrated into the actor network within the MAPPO architecture. For detailed algorithmic implementation, please refer to Appendix F.
Unseen Drone Zoo. To efficiently evaluate adaptive teaming performance, we construct an unseen drone zoo consisting of rule-based methods, traditional optimization-based methods, and learning-based approaches. This pool introduces diverse and challenging teammate behaviors to ensure a robust evaluation of AT-MDP framework in multi-drone pursuit tasks, particularly in continuous-action settings. Specifically, the unseen drone zoo includes three types of algorithms to evaluate adaptive teaming performance. The Greedy Drone follows a rule-based approach, pursuing the closest target while dynamically avoiding obstacles and teammates. The VICSEK Drone, based on the VICSEK algorithm (Janosov et al., 2017a; Zhang et al., 2023; Li et al., 2024a), adopts a bioinspired strategy to mimic swarm-like behaviors, optimizing its path toward the evader while applying repulsive forces to avoid collisions. Finally, the Self-Play Drones use an IPPO-based self-play framework, generating diverse behaviors through randomized training, introducing complexity and variability into the evaluation process. For detailed algorithmic description and implementation, please refer to the Appendix C.
3.2 Simulation and Deployment
Simulation. The simulation module offers a user-friendly and customizable framework through the environment configurator and interface, enabling seamless configuration of multi-drone pursuit scenarios. Fig. 2 (bottom left) shows three examples of custom environments with photos from our real deployed system.
The environment configurator enables users to define all aspects of the simulation by categorizing configuration parameters into three key sections: players, site, and task, as shown in Fig. 7 in Appendix D. This modular design ensures flexibility for a variety of experimental setups. The players section configures participants, including learners, unseen teammates, and evaders, specifying parameters such as their numbers, velocities, and the inclusion of an unseen drone zoo. The site section allows users to customize the physical environment, including map dimensions and obstacle layouts, enabling diverse experimental terrains. Finally, the task section defines the task-specific parameters that govern the rules and objectives of the pursuit scenario.
Environment interface provides a Gymnasium-based interface for reinforcement learning-based training and evaluation. This interface ensures seamless interaction between the training algorithms and the simulation, enabling efficient development and testing of adaptive teaming strategies. By integrating these detailed configurations, the simulation module supports diverse and customized experimental setups while maintaining efficiency for training. Moreover, it bridges the gap between simulation and real-world multi-drone applications, facilitating a robust pipeline for research and deployment.
Deployment. The AT-MDP framework provides a robust real-world deployment solution, integrating edge computing nodes such as the Nvidia Jetson AGX Orin, the OptiTrack motion capture system, and CrazyFlie drones. As shown on the right of Fig. 2, the deployment workflow combines all components into a unified decision-making and execution pipeline. The learners’ policies and unseen drone partners, sampled from the unseen drone pool in the training module, are deployed on the edge nodes, which serve as inference engines for these policies. In this setup, the Jetson AGX Orin edge computing node acts as the core computational unit, offering compact, energy-efficient, on-site processing capabilities. It enables real-time policy inference to support efficient drone decision-making during pursuit tasks. Currently deployed alongside CrazyFlie drones, its lightweight design holds potential for future integration directly onto drones, enabling fully autonomous and decentralized multi-drone systems.
Observations from the physical environment, including drone positions and movements, are captured by the OptiTrack motion capture system, which provides high-precision localization and tracking. The processed observations are sent to the edge nodes, which supply real-time data to the inference engine. The inference engine interprets the data and generates appropriate actions based on the policies deployed. These actions are then communicated to the drones for real-time execution. The drones provide continuous feedback to the system, facilitating dynamic adjustments and enhancing coordination among agents.
Table 2: Performance comparison across different difficulty levels for adaptive teaming without teammate modeling. The means and standard deviations are calculated over five different random seeds, with each seed undergoing 50 repeated runs.
ENV Metrics Unseen Zoo 1 Unseen Zoo 2 Unseen Zoo 3 Unseen Zoo 4
PBT MAPPO OPT PBT MAPPO OPT PBT MAPPO OPT PBT MAPPO OPT
4p2e3o SUC $\uparrow$ 65.20 74.00 77.67 62.00 70.80 86.00 54.80 65.20 77.20 64.80 74.80 82.80
$\pm$ 17.75 $\pm$ 4.69 $\pm$ 6.81 $\pm$ 12.41 $\pm$ 9.23 $\pm$ 3.46 $\pm$ 11.54 $\pm$ 7.69 $\pm$ 3.03 $\pm$ 15.01 $\pm$ 5.40 $\pm$ 1.79
COL $\downarrow$ 32.40 25.60 22.33 27.60 22.40 13.20 39.20 34.00 22.40 31.60 24.00 16.80
$\pm$ 15.19 $\pm$ 4.34 $\pm$ 6.81 $\pm$ 13.30 $\pm$ 12.60 $\pm$ 4.15 $\pm$ 12.30 $\pm$ 7.35 $\pm$ 3.58 $\pm$ 14.38 $\pm$ 4.90 $\pm$ 2.28
AST $\downarrow$ 313.48 303.21 259.17 435.65 386.32 370.95 380.28 376.14 314.97 380.85 328.96 306.35
$\pm$ 86.38 $\pm$ 51.91 $\pm$ 34.28 $\pm$ 113.16 $\pm$ 29.18 $\pm$ 31.36 $\pm$ 68.19 $\pm$ 34.65 $\pm$ 27.68 $\pm$ 85.25 $\pm$ 34.36 $\pm$ 25.86
REW $\uparrow$ 123.60 132.34 138.28 130.51 136.40 149.89 114.00 121.98 135.71 126.15 133.11 143.55
$\pm$ 22.68 $\pm$ 4.68 $\pm$ 7.95 $\pm$ 8.77 $\pm$ 8.50 $\pm$ 3.28 $\pm$ 10.13 $\pm$ 9.07 $\pm$ 3.46 $\pm$ 13.39 $\pm$ 4.68 $\pm$ 1.29
4p2e1o SUC $\uparrow$ 68.80 69.20 80.80 63.20 64.40 73.20 59.20 53.20 67.60 62.40 63.20 75.60
$\pm$ 5.02 $\pm$ 11.71 $\pm$ 7.56 $\pm$ 15.27 $\pm$ 12.20 $\pm$ 3.03 $\pm$ 8.07 $\pm$ 14.11 $\pm$ 5.18 $\pm$ 3.85 $\pm$ 14.46 $\pm$ 5.55
COL $\downarrow$ 30.80 30.40 19.20 31.20 28.80 23.20 36.80 43.20 32.00 34.80 35.60 23.60
$\pm$ 5.22 $\pm$ 12.03 $\pm$ 7.56 $\pm$ 12.85 $\pm$ 15.53 $\pm$ 1.79 $\pm$ 10.73 $\pm$ 15.59 $\pm$ 5.83 $\pm$ 5.40 $\pm$ 14.52 $\pm$ 6.54
AST $\downarrow$ 352.24 317.55 298.93 399.87 383.30 356.65 446.67 395.92 353.35 360.08 375.41 313.60
$\pm$ 30.08 $\pm$ 25.07 $\pm$ 29.19 $\pm$ 86.90 $\pm$ 75.08 $\pm$ 26.78 $\pm$ 72.36 $\pm$ 51.69 $\pm$ 35.29 $\pm$ 48.52 $\pm$ 22.18 $\pm$ 34.18
REW $\uparrow$ 122.05 126.73 138.24 125.85 127.30 134.26 112.66 109.06 120.83 118.73 120.00 131.50
$\pm$ 6.91 $\pm$ 9.92 $\pm$ 6.28 $\pm$ 13.38 $\pm$ 10.62 $\pm$ 3.04 $\pm$ 13.42 $\pm$ 11.94 $\pm$ 6.84 $\pm$ 5.91 $\pm$ 13.22 $\pm$ 5.76
4p2e5o SUC $\uparrow$ 40.40 47.00 50.00 62.40 72.00 75.20 67.60 65.50 74.00 54.80 58.50 67.60
$\pm$ 2.61 $\pm$ 10.39 $\pm$ 11.64 $\pm$ 3.29 $\pm$ 3.65 $\pm$ 6.72 $\pm$ 14.59 $\pm$ 3.42 $\pm$ 4.24 $\pm$ 2.28 $\pm$ 5.97 $\pm$ 6.39
COL $\downarrow$ 59.60 53.00 49.86 35.20 24.00 22.80 31.20 33.50 25.20 44.00 40.50 32.00
$\pm$ 2.61 $\pm$ 10.39 $\pm$ 11.62 $\pm$ 3.90 $\pm$ 5.89 $\pm$ 7.43 $\pm$ 14.74 $\pm$ 3.42 $\pm$ 3.35 $\pm$ 2.45 $\pm$ 5.74 $\pm$ 6.16
AST $\downarrow$ 333.96 313.53 331.98 349.48 419.21 396.66 287.18 348.23 281.45 294.41 340.94 313.79
$\pm$ 34.83 $\pm$ 51.81 $\pm$ 80.30 $\pm$ 17.64 $\pm$ 50.81 $\pm$ 21.48 $\pm$ 18.91 $\pm$ 55.20 $\pm$ 40.77 $\pm$ 38.79 $\pm$ 34.95 $\pm$ 26.78
REW $\uparrow$ 90.16 93.16 99.31 116.19 128.45 134.02 120.48 124.92 128.45 107.56 107.08 119.93
$\pm$ 7.47 $\pm$ 12.00 $\pm$ 17.70 $\pm$ 8.28 $\pm$ 4.27 $\pm$ 6.12 $\pm$ 19.11 $\pm$ 8.33 $\pm$ 3.79 $\pm$ 5.92 $\pm$ 5.44 $\pm$ 6.11
4p3e5o SUC $\uparrow$ 34.80 34.80 31.14 50.40 59.20 67.20 57.20 54.80 60.40 42.00 54.80 60.00
$\pm$ 5.22 $\pm$ 5.40 $\pm$ 17.20 $\pm$ 7.27 $\pm$ 5.40 $\pm$ 5.76 $\pm$ 6.57 $\pm$ 7.43 $\pm$ 5.18 $\pm$ 6.78 $\pm$ 9.44 $\pm$ 6.00
COL $\downarrow$ 62.80 64.80 67.57 46.80 36.00 28.80 40.40 38.00 36.40 55.60 41.60 38.40
$\pm$ 4.60 $\pm$ 5.02 $\pm$ 15.85 $\pm$ 7.56 $\pm$ 4.69 $\pm$ 6.42 $\pm$ 8.05 $\pm$ 5.83 $\pm$ 4.34 $\pm$ 4.77 $\pm$ 8.65 $\pm$ 5.18
AST $\downarrow$ 431.44 509.56 459.55 457.68 489.79 394.81 446.07 555.10 407.85 425.94 510.96 416.91
$\pm$ 67.14 $\pm$ 11.17 $\pm$ 91.71 $\pm$ 33.64 $\pm$ 51.33 $\pm$ 48.65 $\pm$ 38.97 $\pm$ 49.11 $\pm$ 63.06 $\pm$ 60.10 $\pm$ 38.53 $\pm$ 76.21
REW $\uparrow$ 141.98 131.29 116.57 168.82 187.30 193.78 182.04 187.32 196.07 146.74 172.06 174.59
$\pm$ 9.66 $\pm$ 15.98 $\pm$ 43.94 $\pm$ 10.17 $\pm$ 9.08 $\pm$ 4.05 $\pm$ 7.17 $\pm$ 8.97 $\pm$ 9.91 $\pm$ 11.93 $\pm$ 17.01 $\pm$ 18.68
4 Experiment
In this section, we evaluate the performance of baseline methods from the algorithm zoo and validate their feasibility in real-world multi-drone pursuit scenarios. The experiments are structured into two main parts: (1) adaptive teaming without teammate modeling, detailed in Section 4.1, and (2) adaptive teaming with teammate modeling, discussed in Section 4.2. Each subsection presents the experimental setup, key results, and in-depth analysis. Additionally, Section 4.3 provides a case study demonstrating real-world deployment, showcasing how adaptive learners coordinate with unseen drone partners to execute a multi-stage capture strategy. Further details on the real-world implementation are available in Appendix E.
We begin by introducing the shared experimental setups applicable to both groups of experiments.
Environments. To systematically evaluate the performance of adaptive teaming, we leverage the environment configurator to design four multi-drone pursuit environments with varying levels of difficulty, denoted as 4p2e3o, 4p2e1o, 4p2e5o, and 4p3e5o. Fig. 6 in the Appendix gives the screenshots of the four environments deployed in real world. The notation represents the number of pursuers (p), evaders (e), and obstacles (o) in each setting. Each environment is initialized with defined spawn areas, where the evaders spawn within a 3.2m wide and 0.6m high region, while the pursuers spawn in a similarly sized region. The obstacle layouts introduce additional complexity. Environment 4p2e3o include three distributed barriers with different shapes: two cubes in the left area and one cylinder in the right area. Although Environment 4p2e1o only includes a cuboid obstacle in the middle, the difficulty is slight higher than 4p2e3o for evaders have large space to escape. In general, the difficulity of the two environments is easy. Those environments with 5o feature densely packed obstacles that significantly constrain movement. In terms of difficulty, 4p2e3o is categorized as easy due to fewer obstacles and ample space for pursuit, 4p2e1o presents a moderate challenge with a single strategically placed obstacle, 4p2e5o is considered hard as it requires advanced maneuvering in a confined space, and 4p3e5o is the most difficult (superhard) due to an additional evader and dense obstacle layout, requiring highly coordinated teamwork for successful pursuit.
Evaluation Metrics. To evaluate the performance of the methods, we use the following metrics: 1. Success rate (SUC), which measures the proportion of tasks successfully completed. An episode is deemed successful if the pursuers capture both evaders, defined as reducing the distance between an evader and a pursuer to less than 0.2 metres; 2. Collision rate (COL), which tracks the frequency of collisions during task execution. A collision is recorded if the distance between any two pursuers is less than 0.2 metres or if the distance between a drone and an obstacle is less than 0.1 metres; 3. Average success timesteps (AST), which indicates the average number of steps taken to complete a task; 4. Average reward (REW), which reflects the overall efficiency and quality of the agent’s performance across episodes. These metrics collectively provide a comprehensive evaluation of the proposed method’s adaptability and effectiveness.
4.1 Adaptive teaming without teammate modeling
4.1.1 Evaluation Protocol
To ensure a rigorous and fair comparison of adaptive teaming, we standardize the evaluation process across all methods. Since AT-MDP is a newly introduced problem with no existing algorithms specifically designed for it, we establish baselines using PBT (Carroll et al., 2019), MAPPO (Yu et al., 2022), and our proposed OPT.
Unseen Drone Partners. Unseen drone partners are a critical component of the evaluation protocol. To thoroughly assess the proposed method, we evaluate its performance across four distinct unseen drone zoos, with drone policies selected from the unseen drone pool in the AT-MDP framework. Each zoo represents a unique combination of drone behaviors, designed to create diverse and challenging scenarios that test the adaptability and robustness of the proposed method. (1) Unseen Zoo 1 and (2) Unseen Zoo 3 include a Greedy drone policy and a VICSEK drone policy (Janosov et al., 2017a), respectively, providing additional diversity in drone behavior and coordination strategies. (3) Unseen Zoo 3 consists of two drone policies trained using a self-play PPO algorithm (Schulman et al., 2017; Carroll et al., 2019) with different random seeds. These policies are selected to reflect varying levels of coordination ability: one with strong coordination skills and another with relatively weaker performance. For example, in a 4p2e1o scenario, the success rates (SUC) of the two policies, when evaluated against themselves, are 70% and 54%, respectively. (4) Unseen Zoo 4 combines all policies from the other three zoos into a single zoo. At the beginning of each episode, unseen drone partners are randomly selected from this combined pool, introducing additional variability and unpredictability to the evaluation process.
4.1.2 Main Results
Performance Across Different Environments. As the complexity of the environments increases (from 4p2e3o to 4p3e5o), the success rate (SUC) generally decreases across all methods, highlighting the increasing difficulty in coordinating with unseen teammates. In simpler environments like 4p2e3o and 4p2e1o, OPT consistently outperforms other baselines, achieving higher success rates and lower collision rates. In more challenging settings (4p2e5o and 4p3e5o), performance gaps become more pronounced, with MAPPO and PBT struggling to generalise effectively.
Effectiveness of OPT in Adaptive Teaming. OPT achieves the highest success rates in most cases, particularly in Unseen Zoo 2 and Unseen Zoo 4, demonstrating its robustness in adapting to diverse teammate behaviors. It significantly reduces AST, meaning it captures evaders faster than other baselines, suggesting that its learning process leads to more efficient pursuit strategies. Moreover, OPT consistently achieves higher reward values, indicating its superior adaptability in unseen teaming scenarios.
Impact of Unseen Drone Zoos. The performance of all methods varies significantly across different unseen drone zoos, each introducing distinct teammate behaviors. Unseen zoo 1 (Greedy agents) and Unseen zoo 2 (VICSEK-based agents) exhibit more predictable behaviors, allowing for more effective coordination, with all methods achieving relatively higher SUC scores. Unseen zoo 3, composed entirely of PPO-based self-play policies with two different skill levels, presents the most challenging setting, where all baselines experience performance drops due to the increased adaptability and unpredictability of teammates. Unseen zoo 4, which includes a mix of all drone types with random selection, further increases variability, testing the generalization ability of each method. Notably, OPT remains the most stable approach, demonstrating robust adaptability across diverse teammate dynamics.
4.2 Adaptive teaming with teammate modeling
We also conduct a group of experiments about adatptive teaming with teammates modeling. As shown Table. 3 presents the performance comparison between MAPPO and ATM across different difficulty levels. As the environment complexity increases (from 4p2e3o to 4p3e5o), both methods exhibit a decline in success rate (SUC) and an increase in average steps to capture (AST), highlighting the growing challenge of coordinating with teammates in more constrained and adversarial settings. ATM consistently outperforms MAPPO in success rate across all environments, demonstrating its improved adaptability in dynamic multi-agent interactions. Notably, ATM achieves a significantly higher success rate in 4p2e3o and 4p2e1o, suggesting its effectiveness in simpler settings. However, as the complexity increases (4p2e5o and 4p3e5o), the performance gap narrows, indicating the growing challenge of maintaining robust coordination under more difficult conditions. In terms of collision rate (COL), ATM generally results in lower or comparable values to MAPPO, suggesting that its teammate modeling mechanism helps mitigate unnecessary collisions. The reward metric (REW) further supports these findings, as ATM achieves higher values across all settings, reflecting its superior learning of adaptive strategies.
4.3 Case Study
To further illustrate the effectiveness of our adaptive teaming approach, we present a case study in the 4p3e5o environment, categorized as superhard due to its high complexity, featuring four pursuers, three evaders, and five obstacles. The unseen teammates in this scenario are sampled from Unseen Zoo 3, which consists entirely of PPO-based self-play policies trained at two different skill levels, introducing high adaptability and unpredictability.
This case study demonstrates how the ATM learners effectively coordinate with their unseen drone partners to execute a multi-stage capture strategy. Fig. 3 illustrates key frames from the scenario. In Frame 1, four pursuers initiate a collaborative approach, positioning themselves strategically to encircle all three evaders while maintaining an adaptive formation. In Frame 2, two pursuers successfully capture one of the evaders while the other two tighten their formation, preventing the remaining evaders from escaping. In Frames 3 & 4, the remaining two evaders are captured one by one as the pursuers continue refining their positioning and coordination, effectively closing all escape routes.
Table 3: Performance comparison across different difficulty levels. Metrics are reported as mean $\pm$ standard deviation.
Metrics 4p2e3o 4p2e1o 4p2e5o 4p3e5o
MAPPO ATM MAPPO ATM MAPPO ATM MAPPO ATM
SUC $\uparrow$ 72.80 77.20 71.60 72.00 52.40 72.40 51.20 62.00
($\pm$ 4.15) ($\pm$ 6.87) ($\pm$ 3.29) ($\pm$ 4.00) ($\pm$ 5.55) ($\pm$ 5.55) ($\pm$ 7.95) ($\pm$ 8.00)
COL $\downarrow$ 26.40 22.00 26.40 26.40 47.60 26.80 48.00 38.00
($\pm$ 4.56) ($\pm$ 7.62) ($\pm$ 3.58) ($\pm$ 6.07) ($\pm$ 5.55) ($\pm$ 6.10) ($\pm$ 7.48) ($\pm$ 8.00)
AST $\downarrow$ 297.46 277.55 370.15 326.03 289.01 283.59 408.28 372.33
($\pm$ 37.98) ($\pm$ 17.55) ($\pm$ 65.15) ($\pm$ 48.37) ($\pm$ 26.45) ($\pm$ 45.80) (±($\pm$( ± 45.57) ($\pm$ 17.39)
REW $\uparrow$ 133.03 136.84 128.79 130.08 104.79 128.45 164.19 180.54
($\pm$ 4.21) ($\pm$ 6.91) ($\pm$ 1.79) ($\pm$ 4.71) ($\pm$ 3.76) ($\pm$ 6.17) ($\pm$ 14.67) ($\pm$ 15.67)
Figure 3: Case Study: This example demonstrates the capture strategy executed by ATM learners and unseen drone partners from unseen zoo 3 in the superhard environment 4p3e5o. The red circles denote pursuers, and the black squares represent evaders. In this scenario, four pursuers collaboratively surround all three evader (1), two pursuers capture one of evaders while other two pursuers continuously tighten their formation (2), and rest of two evaders are then successfully captured one by one (3 & 4)
5 Conclusion
In this paper, we define and formalize the adaptive teaming in the multi-drone pursuit problem and introduce AT-MDP framework, a comprehensive framework that integrates simulation, training, and real-world deployment to address this challenge. AT-MDP framework provides a user-friendly and flexible environment configurator, a distributed training framework incorporating an extensive algorithm zoo and an unseen drone zoo, and a real-world deployment system leveraging edge computing and Crazyflie drones. To the best of our knowledge, AT-MDP framework is the first adaptive teaming framework designed for continuous-action multi-drone pursuit tasks, extending multi-agent coordination beyond pre-coordinated strategies. Our experiments in four environments of increasing difficulty validate the effectiveness of our proposed baseline methods, demonstrating their adaptability to unseen teammates and superior performance compared to existing approaches. Furthermore, real-world experiments confirm the feasibility of AT-MDP framework in physical drone systems, bridging the gap between simulation and real-world deployment.
Limitations and Future Work. While AT-MDP framework successfully bridges simulation and real-world deployment, the current real-world system is relatively simple, relying on Crazyflie drones without onboard perception, LiDAR, or computer vision capabilities. Furthermore, the scenarios tested, while diverse, could be further enhanced to better reflect real-world complexities. Future work will focus on integrating perception modules, expanding the framework to more realistic environments, and developing a multi-drone system equipped with edge computing devices, such as Jetson, to enable fully autonomous adaptive teaming.
6 Impact Statement
This paper advances the field of machine learning and multi-robot systems by introducing Adaptive Teaming in Multi-Drone Pursuit (AT-MDP) and proposing AT-MDP framework, a comprehensive framework for simulation, training, and real-world deployment. Our work contributes to multi-drone system by enabling multiple drones to dynamically coordinate with unseen teammates in continuous-action real-world tasks.
The potential societal benefits of this research include improvements in search-and-rescue operations, disaster response, border surveillance, and autonomous security systems, where adaptive coordination among unseen agents can enhance efficiency and safety. Additionally, the open-source implementation of our framework fosters reproducibility and further research in adaptive teaming and multi-agent collaboration.
From an ethical perspective, while multi-drone pursuit has positive real-world applications, the technology also has potential risks if misused, such as surveillance or adversarial applications. We emphasize the importance of responsible AI development and adherence to ethical guidelines in deploying autonomous systems. Future work should consider human-in-the-loop oversight and safety constraints to ensure the ethical use of AI-driven multi-agent systems.
Appendix A Related Work
Table 4: Comparison of AT-MDP framework with related works. Grey rows represent literature related to multi-drone pursuit, while pink rows highlight adaptive teaming studies from the machine learning field. “AT w/o TM” denotes adaptive teaming without teammate modelling, while “AT w/ TM” refers to adaptive teaming with teammate modelling.
Related Work Problem Setting Task Method
# Learner # Unseen # Evader Action Space Main Related Task Real-world? AT w/o TM? AT w/ TM?
Voronoi Partitions (Zhou et al., 2016) Multi 0 1 Continuous Pursuit–evasion Game No No No
Bio-pursuit (Janosov et al., 2017b) Multi 0 Multi Continuous Prey–predator Game No No No
Uncertainty-pursuit (Shah & Schwager, 2019) Multi 0 1 Continuous Pursuit–evasion Game No No No
M3DDPG (Li et al., 2019) Multi 0 1 Continuous Prey–predator Game No No No
Pursuit-TD3(de Souza et al., 2021) Multi 0 1 Continuous Multi-drone Pursuit Yes No No
DACOOP-A(Zhang et al., 2023) Multi 0 1 Discrete Multi-drone Pursuit Yes No No
GM-TD3 (Zhang et al., 2024) Multi 0 1 Continuous Prey–predator Game No No No
DualCL (Chen et al., 2024) Multi 0 1 Continuous Multi-drone Pursuit No No No
HOLA-Drone (Li et al., 2024a) 1 Multi Multi Continuous Multi-drone Pursuit Yes Yes No
Other-play (Hu et al., 2020) 1 1 0 Discrete Lever Game; Hanabi No Yes No
Overcooked-AI  (Carroll et al., 2019) 1 1 0 Discrete Overcooked No Yes No
TrajDi (Lupu et al., 2021) 1 1 0 Discrete Overcooked No Yes No
MEP (Zhao et al., 2023) 1 1 0 Discrete Overcooked No Yes No
LIPO (Charakorn et al., 2023) 1 1 0 Discrete Overcooked No Yes No
COLE (Li et al., 2024b) 1 1 0 Discrete Overcooked No Yes No
ZSC-Eval (Wang et al., 2024d) 1 1 0 Discrete Overcooked No Yes No
PLASTIC (Barrett et al., 2017) 1 Multi Multi Discrete Prey-predator Game No No Yes
AATeam (Chen et al., 2020) 1 1 2 Discrete Half Field Offense No No Yes
LIAM (Papoudakis et al., 2021) 1 Multi Multi Discrete LBF; Prey-predator Game No No Yes
GPL (Rahman et al., 2021) 1 Multi Multi Discrete LBF; Wolfpack; FortAttack No No Yes
CIAO (Wang et al., 2024b) 1 Multi Multi Discrete LBF; Wolfpack No No Yes
NAHT (Wang et al., 2024a) Multi Multi Multi Discrete StarCraft2 No No Yes
Our ATMDP Multi Multi Multi Continuous Multi-drone Pursuit Yes Yes Yes
In this work, we provide a comprehensive review of related research on multi-drone pursuit and adaptive teaming in machine learning, with a detailed comparison presented in Table 1.
Multi-agent pursuit-evasion. Multi-agent pursuit-evasion is closely related to the multi-drone pursuit task. Most existing methods rely on pre-coordinated strategies specifically designed for particular pursuit-evasion scenarios. Traditional approaches often rely on heuristic (Janosov et al., 2017b) or optimisation-based strategies (Zhou et al., 2016; Shah & Schwager, 2019). For example, Janosov et al. (2017b) proposes a bio-inspired model that uses local interaction rules to enhance group chasing success in Prey–Predator Games. Similarly, the Voronoi partitions algorithm (Zhou et al., 2016) and the uncertainty-pursuit algorithm (Shah & Schwager, 2019) employ decentralised frameworks to optimise the evader’s Voronoi partition and reachable area, respectively. In recent years, deep reinforcement learning (DRL) has been widely adopted for pre-coordinated multi-drone pursuit tasks. M3DDPG (Li et al., 2019) and GM-TD3 (Zhang et al., 2024) extend standard DRL algorithms, such as TD3 (Fujimoto et al., 2018) and DDPG (Lillicrap, 2015), specifically for multi-agent pursuit in simulated environments. Pursuit-TD3 (de Souza et al., 2021) applies the TD3 algorithm to pursue a target with multiple homogeneous agents, validated through both simulations and real-world drone demonstrations. Zhang et al. (2023) introduces DACOOP-A, a cooperative pursuit algorithm that enhances reinforcement learning with artificial potential fields and attention mechanisms, validated in real-world drone systems. DualCL (Chen et al., 2024) addresses multi-UAV pursuit-evasion in diverse environments and demonstrates zero-shot transfer capabilities to unseen scenarios, though only in simulation. The most recent work, HOLA-Drone (Li et al., 2024a), claims to be the first zero-shot coordination framework for multi-drone pursuit. However, it is limited to controlling a single learner, restricting its applicability to broader multi-agent settings.
Adaptive Teaming. The adaptive teaming paradigm can be broadly categorised into two approaches: adaptive teaming without teammate modelling (AT w/o TM) and adaptive teaming with teammate modelling (AT w/ TM), which correspond to the zero-shot coordination (ZSC) and ad-hoc teamwork (AHT) problems in the machine learning community, respectively. AT w/o TM focuses on enabling agents to coordinate with unseen teammates without explicitly modelling their behaviours. Other-Play (Hu et al., 2020) introduces an approach that leverages symmetries in the environment to train robust coordination policies, applied to discrete-action tasks like the Lever Game and Hanabi. Similarly, methods such as Overcooked-AI (Carroll et al., 2019), TrajDi (Lupu et al., 2021), MEP (Zhao et al., 2023), LIPO (Charakorn et al., 2023), and ZSC-Eval (Wang et al., 2024d) study collaborative behaviours in Overcooked, where agents learn generalisable coordination strategies with diverse unseen partners. While these approaches demonstrate promising results, they are limited to single-learner frameworks in simplified, discrete-action domains like Overcooked and Hanabi. They lack scalability to multi-agent settings, continuous action spaces, and the complexities of real-world applications.
AT w/ TM, on the other hand, explicitly models the behaviour of unseen teammates to facilitate effective collaboration. Early methods like PLASTIC (Barrett et al., 2017) reuse knowledge from previous teammates or expert input to adapt to new teammates efficiently. AaTeam (Chen et al., 2020) introduces attention-based neural networks to dynamically process and respond to teammates’ behaviours in real-time. More advanced approaches, such as LIAM (Papoudakis et al., 2021), employ encoder-decoder architectures to model teammates using local information from the controlled agent. GPL (Rahman et al., 2021) and CIAO (Wang et al., 2024b) leverage graph neural networks (GNNs) to address the challenges of dynamic team sizes in AHT. Extending this further, NAHT (Wang et al., 2024a) enables multiple learners to collaborate and interact with diverse unseen partners in N-agent scenarios. Despite their progress, these methods remain confined to discrete action spaces and simulated benchmarks, limiting their applicability to real-world, continuous-action tasks.
Figure 4: Overview of our proposed open-ended population training algorithm.
Appendix B Open-ended Population Training Algorithm
In this section, we define a population of drone strategies, denoted as $\Pi=\{\pi_{1},\pi_{2},\cdots,\pi_{n}\}$. For the task involving $C$ teammates, the interactions within the population $\Pi$ are modeled as a hypergraph $\mathcal{G}$. Formally, the hypergraph is represented by the tuple $(\Pi,\mathcal{E},\mathbf{w})$, where the node set $\Pi$ represents the strategies, $\mathcal{E}$ is the hyperedge set capturing interaction relationships among teammates, and $\mathbf{w}$ is the weight set representing the corresponding average outcomes. The left subfigure of Fig. 5 illustrates an example of a hypergraph representation with five nodes and a fixed hyperedge length of 4.
Building on the concept of preference hypergraphs (Li et al., 2024a), we use the preference hypergraph to represent the population and assess the coordination ability of each node. The preference hypergraph $\mathcal{P}\mathcal{G}$ is derived from the hypergraph $\mathcal{G}$, where each node has a direct outgoing hyperedge pointing to the teammates with whom it achieves the highest weight in $\mathcal{G}$. Formally, $\mathcal{P}\mathcal{G}$ is defined by the tuple $(\Pi,\mathcal{E}_{\mathcal{P}})$, where the node set $\Pi$ represents the strategies, and $\mathcal{E}_{\mathcal{P}}$ denotes the set of outgoing hyperedges. As shown in the right subfigure of Fig. 5, the dotted line highlights the outgoing edge. For instance, node $2$ has a single outgoing edge $(2,3,5,4)$ because it achieves the highest outcome, i.e., a weight of 45, with those teammates in $\mathcal{G}$, as depicted in the left subfigure.
Intuitively, a node in $\mathcal{P}\mathcal{G}$ with higher cooperative ability will have more incoming hyperedges, as other agents prefer collaborating with it to achieve the highest outcomes. Therefore, we extend the concept of preference centrality (Li et al., 2023) to quantify the cooperative ability of each node. Specifically, for any node $i\in\Pi$, the preference centrality is defined as
$\eta_{\Pi}(i)=\frac{d_{\mathcal{P}\mathcal{G}}(i)}{d_{\mathcal{G}}(i)},$ (1)
where $d_{\mathcal{P}\mathcal{G}}(i)$ denotes the incoming degree of node $i$ in $\mathcal{P}\mathcal{G}$, and $d_{\mathcal{G}}(i)$ represents the degree of node $i$ in $\mathcal{G}$.
Max-Min Preference Oracle. Building on the basic definition of the preference hypergraph representation, we introduce the concept of Preference Optimality to describe the goal of our training process.
Figure 5: An example of a hypergraph representation (left) and its corresponding preference hypergraph (right) with five strategies in the population.
Definition B.1 (Preference Optimal).
A set of learners $\mathcal{N}^{\star}$ of size $N$ is said to be Preference Optimal (PO) in a hypergraph $\mathcal{G}=(\Pi,\mathcal{E},\mathbf{w})$ if, for any set $\hat{\mathcal{N}}\subseteq\Pi$ of size $N$, the following condition holds:
$\sum_{s\in\mathcal{N}^{\star}}\eta_{\Pi}(s)\geq\sum_{s\in\hat{\mathcal{N}}}\eta_{\Pi}(s),$ (2)
where $\eta_{\Pi}(s)$ denotes the preference centrality of learner $s$ in the hypergraph $\mathcal{G}$.
While achieving a preference-optimal oracle is desirable, it becomes impractical or prohibitively expensive in large, diverse populations. Therefore, we propose the max-min preference oracle, abbreviated as oracle in the rest of this paper, to ensure robust adaptability and maximize cooperative performance under the worst-case teammate scenarios.
To formalize the objective, we split the strategy population $\Pi$ into a learner set $\mathcal{N}$ and a non-learner set $\Pi_{-\mathcal{N}}$, where $\Pi_{-\mathcal{N}}\cap\mathcal{N}=\emptyset$ and $\Pi_{-\mathcal{N}}\cup\mathcal{N}=\Pi$. The objective function $\phi$ is defined as:
$\phi:\underbrace{\mathcal{N}\times\cdots\times\mathcal{N}}_{N\text{ learners}}\times\underbrace{\Pi_{-\mathcal{N}}\times\cdots\times\Pi_{-\mathcal{N}}}_{M\text{ teammates}}\rightarrow\mathbb{R}.$ (3)
The max-min preference oracle updates the learner set by solving:
$\mathcal{N}^{\prime}$ $=oracle(\mathcal{N},\phi_{\mathcal{M}}(\cdot)):=\arg\max_{\mathcal{N}}\min_{\mathcal{M}\subseteq\Pi_{-\mathcal{N}}}\phi_{\mathcal{M}}(\mathcal{N}),$ (4)
where the objective $\phi_{\mathcal{M}}(\cdot)$ is derived using the extended curry operator (Balduzzi et al., 2019), originally designed for two-player games, and is expressed as:
$\left[\underbrace{\mathcal{N}\times\cdots\times\mathcal{N}}_{N\text{ learners}}\times\underbrace{\Pi_{-\mathcal{N}}\times\cdots\times\Pi_{-\mathcal{N}}}_{M\text{ teammates}}\rightarrow\mathbb{R}\right]$ (5)
$\rightarrow$ $\left[\underbrace{\Pi_{-\mathcal{N}}\times\cdots\times\Pi_{-\mathcal{N}}}_{M\text{ teammates}}\rightarrow\left[\underbrace{\mathcal{N}\times\cdots\times\mathcal{N}}_{N\text{ learners}}\rightarrow\mathbb{R}\right]\right].$
Intuitively, oracle alternates between two key steps: the minimization step and the maximization step. In the minimization step, the objective is to identify the subset of teammates $\mathcal{M}^{*}\subset\Pi_{-{\mathcal{N}}}$ that minimizes the performance outcome of the current learner set $\mathcal{N}$, i.e. the worst partners. This is formulated as:
$\mathcal{M}^{*}=\arg\min_{\mathcal{M}\subset\Pi_{-{\mathcal{N}}}}\phi_{\mathcal{M}}(\mathcal{N}).$
In the maximization step, the learner set $\mathcal{N}$ is updated to maximize its performance outcome against the identified subset $\mathcal{M}^{*}$. This is defined as:
$\mathcal{N}^{*}=\arg\max_{\mathcal{N}}\phi(\mathcal{N},\mathcal{M}^{*}).$
Open-Ended Population Training algorithm. To achieve robust adaptability and dynamic coordination in multi-agent systems, we integrate the max-min preference oracle into an open-ended learning framework, referred to as the Open-ended Population Training (OPT) algorithm. The OPT algorithm dynamically adjusts the training objective as the population evolves, enabling continuous improvement and effective coordination with unseen partners. Unlike conventional fixed-objective training, the OPT approach iteratively expands the strategy population $\Pi$ and refines the learner set $\mathcal{N}$. At each generation $t$, the framework recalibrates the training objective $\phi$ based on new extended population $\Pi_{t}$ to account for the evolving interactions among agents within the population.
As shown in Fig. 5, the Open-Ended Population Training (OPT) algorithm consists of two key modules: the min-step solver and the max-step trainer. At each generation $t$, the updated learner set $\mathcal{N}_{t}$ from the previous generation $t-1$ is incorporated into the population $\Pi_{t-1}$, resulting in an expanded population $\Pi_{t}$.
Min-step Solver. The role of the min-step solver is to first construct the preference hypergraph representation of the interactions within the updated population $\Pi_{t}$. Here, we only need to build a subgraph of the entire preference hypergraph in $\Pi_{t}$, denoted as $\mathcal{P}\mathcal{G}_{t}^{\prime}$. To obtain $\mathcal{P}\mathcal{G}_{t}^{\prime}$, we focus on constructing the hyperedges in the hypergraph $\mathcal{G}_{t}^{\prime}$ that connect to the learner set $\mathcal{N}_{t}$. For instance, if $\Pi_{t}$ consists of a learner set $\mathcal{N}_{t}$ of size $N$ and a non-learner set $\Pi_{-\mathcal{N}_{t}}$, any hyperedge $e$ in $\mathcal{G}_{t}^{\prime}$ connects $N$ nodes from $\mathcal{N}_{t}$ and all possible $M$ nodes from $\Pi_{-\mathcal{N}_{t}}$. The preference hypergraph $\mathcal{P}\mathcal{G}_{t}^{\prime}$ is then derived from $\mathcal{G}_{t}^{\prime}$ by retaining only the outgoing hyperedge with the highest weight for each node.
The min-step solver uses the reciprocal of the preference centrality in $\mathcal{P}\mathcal{G}_{t}^{\prime}$ to evaluate the worst-case partners. To enhance robustness, the solver does not deterministically select the worst-case partners as $\mathcal{M}^{*}=\arg\min_{\mathcal{M}\subset\Pi_{-\mathcal{N}}}\phi_{\mathcal{M}}(\mathcal{N})$. Instead, it outputs a mixed strategy $\rho_{t}$, defined as:
$\rho_{t}=\arg\min_{P(\Pi_{-\mathcal{N}_{t}})}\mathbb{E}_{\mathcal{M}\sim P}[\phi_{\mathcal{M}}(\mathcal{N}_{t})].$ (6)
In practice, the mixed strategy $\rho_{t}$ is obtained by normalizing the reciprocal of the preference centrality, assigning higher probabilities to worse partners.
Max-Step Solver. Given the mixed strategies $\rho_{t}$, the max-step solver iteratively samples the worst-case partners, referred to as the profile, $\mathcal{M}\sim\rho_{t}$, from the non-learner set $\Pi_{-\mathcal{N}_{t}}$. It simulates interactions between the sampled profile and the learners to generate training data, with the objective of maximizing the reward $\phi(\mathcal{N}_{t},\mathcal{M})=\mathbb{E}_{\mathcal{N}_{t},\mathcal{M}}[R(\tau)]$, as shown in Eq. 2.1. The max-step oracle could be rewritten as
$\mathcal{N}_{t+1}=\arg\max_{\mathcal{N}}\mathbb{E}_{\mathcal{M}^{\star}\sim\rho_{t}}\phi(\mathcal{N},\mathcal{M}^{*}).$ (7)
Figure 6: Illustration of four multi-drone pursuit environments in real world. The environments vary in the number of pursuers (p), evaders (e), and obstacles (o), denoted as 4p2e3o, 4p2e1o, 4p2e5o, and 4p3e5o. Each setup introduces different levels of complexity, testing the adaptability and coordination capabilities of the agents.
The strategy network for adaptive teaming, supports both AHT and ZSC paradigms. In the AHT paradigm, the network uses a teammate modeling network ($f$) to infer teammate types from the observation history ($\tau_{t}^{i}$). These predicted vector, combined with the agent’s observation history ($\tau_{t}^{i}$), are input into a PPO-based policy network. This policy network includes an Actor Network ($\pi_{\theta}$) for generating the agent’s action ($a_{t}^{i}$) and a Critic Network ($V_{\pi}$) for evaluating the policy.
In contrast, the ZSC paradigm simplifies the process by directly feeding the agent’s observation history ($\tau_{t}^{i}$) into the actor and critic networks, bypassing explicit teammate modeling. This approach enables the agent to coordinate with unseen teammates without prior knowledge or additional inference mechanisms.
The max-step solver ultimately generates an approximate best response ${\mathcal{N}}_{t+1}$ to the worst-case partners, enhancing the agent’s adaptive coordination capabilities.
Appendix C Unseen Drone Zoo
Rule-Based Method: Greedy Drone. The Greedy Drone pursues the closest target by continuously aligning its movement with the target’s position. Its state information includes its own position, orientation, distances and angles to teammates and evaders, and proximity to obstacles or walls. When obstacles or other agents enter its evasion range, the Greedy Drone dynamically adjusts its direction to avoid collisions, prioritising immediate objectives over team coordination.
Traditional Method: VICSEK Drone. Based on the commonly used VICSEK algorithm (Janosov et al., 2017a; Zhang et al., 2023; Li et al., 2024a), the VICSEK Drone adopts a bio-inspired approach to mimic swarm-like behaviours. It computes and updates a velocity vector directed towards the evader, optimising the tracking path based on the agent’s current environmental state. To avoid nearby obstacles or agents, the VICSEK Drone applies repulsive forces with varying magnitudes. While the calculated velocity vector includes both magnitude and orientation, only the orientation is implemented in our experiments, making it a scalable and practical teammate model for multi-drone coordination.
Learning-Based Method: Self-Play Drones. For the learning-based approach, we employ an IPPO-based self-play algorithm, generating diverse drone behaviours by training agents with different random seeds. This approach simulates a wide range of adaptive strategies, introducing stochasticity and complexity to the evaluation process.
"players": {
"num_p": 4,
"num_e": 2,
"num_ctrl": 2,
"num_unctrl": 2,
"random_respawn": True,
"respawn_region": {***},
"reception_range": 2,
"velocity_p": 0.3,
"velocity_e": 0.6,
"unseen_drones": [***]
"site":{
"boundary": {
"width" : 3.6,
"height" : 5,
"obstacles": {
"obstacle1":{***}
"task":{
"task_name": 4p2e1o,
"capture_range": 0.2,
"safe_radius": 0.1,
"task_horizon": 100,
"fps": 10,
Figure 7: An example of environment configuration file.
Appendix D Environment Configurator
The AT-MDP framework environment configurator allows users to define and modify multi-drone pursuit scenarios through a structured JSON file. Fig. 7 provides an example configuration file that specifies key parameters across three categories: players, site, and task.
Players Configuration: This section defines the number and roles of agents in the environment, including the number of pursuers (num_p), evaders (num_e), controlled agents (num_ctrl), and unseen teammates (num_unctrl). Additional parameters such as random respawn behavior, reception range, and velocity settings further customize agent interactions. The unseen_drones field allows users to specify different unseen teammate models from the unseen drone zoo.
Site Configuration: This section defines the physical properties of the environment, including its boundary dimensions (width, height) and obstacle placements. Obstacles can be configured individually to introduce varying levels of complexity.
Task Configuration: This section sets the pursuit task parameters, including the capture range (capture_range), safety radius (safe_radius), task duration (task_horizon), and simulation frame rate (fps). The task_name field provides a label for different predefined environment scenarios.
This modular configuration enables flexible environment customization, facilitating experiments across diverse multi-drone pursuit scenarios.
Appendix E Real-world Deployment
To validate the feasibility of our algorithms in real-world scenarios, we design and conduct hardware experiments using Crazyflie drones, the FZMotion motion capture system, and the Crazyswarm validation platform. The FZMotion motion capture system is responsible for real-time position measurement of the drones. It transmits the positional data in point cloud format to Crazyswarm, where the information is processed as input for the algorithm. The Crazyswarm platform is deployed on two different edge nodes, i.e. a laptop ( Lenovo ThinkPad T590) and a Jetson Orin Nano, which handles the reception of drone position data from the motion capture system, executes the adaptive teaming algorithm, and transmits control commands to the Crazyflie drones via Crazyradio PA. Upon receiving these commands, the Crazyflie drones execute the prescribed maneuvers. The onboard Mellinger controller ensures accurate trajectory tracking, allowing the drones to follow the control signals with high precision. This real-world deployment setup enables the direct evaluation of our learned policies in physical drone systems, bridging the gap between simulation and real-world execution.
Table 5: Implementation hyperparameters of ATM baseline algorithm.
Parameters Values Parameters Values
Batch size 1024 Minibatch size 256
Lambda ($\lambda$) 0.99 Generalized advantage estimation lambda ($\lambda_{gae}$) 0.95
Learning rate 3e-4 Value loss coefficient($c_{1}$) 1
Entropy coefficient($\epsilon_{clip}$) 0.01 PPO epoch 20
Total environment step 1e6 History length 1
Embedding size 16 Hidden size 128
Appendix F Implementation Details of Adaptive Teaming with Modeling
Adaptive Teaming with Modeling (ATM) extends the MAPPO algorithm (Yu et al., 2022) by incorporating an additional teammate modeling network $f$. This network generates team encoding vectors to represent the behavioral characteristics of unseen teammates, improving coordination in multi-drone pursuit. The modeling network $f$ processes three types of inputs: (1) observed evader states history, (2) self-observed states history, and (3) relative positions history between agents. These inputs are transformed using independent fully connected layers, aggregated through weighted averaging, and combined into a unified team representation. The final embedding is a fixed-dimensional vector, integrated into the actor network of MAPPO to enhance decision-making in adaptive teaming scenarios. The details of hyperparameters are listed in Table 5.